OpenStudy (anonymous):

solve y=x2-6x+10 using general form ? helppppp pleaseeee asap..

OpenStudy (ganpat):

equate the equation to 0.. x2-6x+10= 0 and solve by factorization.. can u do that ??

OpenStudy (ganpat):

h aha.. ok.. just show her the troll face, she will understand.. kidding.. follow this steps carefully... x2-6x+10 = 0 well, i dont think it can be factorized ... we have to use other formula..

OpenStudy (ganpat):

x2-6x+10 compare it with ax2 + bx + c = 0 and use this formula..